OpenStudy (anonymous):

half-angle formula: I'm trying to integrate the sqrt 2(1+cosx) and I know you can use half-angle to turn it into sqrt 4(cos^2(x)). But I don't see *how* it turned into that :( Can someone please explain

OpenStudy (anonymous):

I think it's the second identity in this link: http://www.freemathhelp.com/images/halfangles.png and I think you made a typo

OpenStudy (jhannybean):

\[\color{red}{\int \sqrt{2(1+\cos(x))}dx}\]\[=\int \sqrt{2} \cdot \sqrt{1+\cos(x)}dx\]\[=\sqrt{2}\int\sqrt{1+\cos(x)}dx\]If \(\cos(2x) =2\cos^2(x)-1\), then \(\cos(x) = 2\cos^2\left(\dfrac{x}{2}\right)-1\implies \color{blue}{\cos(x)+1=2\cos^2\left(\dfrac{x}{2}\right)} \)\[=\sqrt{2}\int \sqrt{2\cos^2\left(\dfrac{x}{2}\right)}dx\]\[=\sqrt{2}\int \sqrt{2}\sqrt{\cos^2\left(\dfrac{x}{2}\right)}dx \]\[=2\int\sqrt{\cos^2\left(\dfrac{x}{2}\right)}dx\]Let \(u=\dfrac{x}{2} ~,~ du = \dfrac{1}{2}dx \implies 2du = dx\)

OpenStudy (jhannybean):

\[=4\int \sqrt{\cos^2(u)}du\]\[4\int \cos(u)du =\color{red}{4\sin(u)+C} = 4\sin\left(\dfrac{x}{2}\right)+C\]Can be simplified one more step but I'll let you figure that out.

OpenStudy (dan815):

\[\int\sqrt {4(cos^2(x/2))} \]

OpenStudy (dan815):

dx

OpenStudy (anonymous):

It's good

OpenStudy (anonymous):

For the second one \[2 \int\limits \sqrt{\cos^2(x/2)}dx = 4 \int\limits \sqrt{\cos^2u}du\] letting u = x/2 and du = 1/2 dx you can finish it off :)

OpenStudy (anonymous):

About 1-2 steps away hehe

OpenStudy (jhannybean):

Water u talkin' bout mate?

OpenStudy (anonymous):

The second integral

OpenStudy (anonymous):

Oh wait, is it suppose to be the same one haha?

OpenStudy (jhannybean):

Yeah the squared and the square root cancel eachother out.

OpenStudy (anonymous):

I was doing what dan posted

OpenStudy (jhannybean):

Yeah it's the same one, hence the equal sign lmao -.-

OpenStudy (anonymous):

Yeah ok that makes sense lol I just realized that, because I just went over your work

OpenStudy (anonymous):

then dan wrote that

OpenStudy (jhannybean):

My post was just lagging really bad and I was afraid it was going to disappear like the first time D::

OpenStudy (anonymous):

:)